# Wrapyfi: A Python Wrapper for Integrating Robots, Sensors, and Applications across Multiple Middleware

### 12 Dec 2023 v4

@@ Abstract.
https://github.com/fabawi/wrapyfi.
@@ 1. Introduction
To improve interoperability between different robotic platforms and reduce reliance on a particular middleware, we have developed the open source Wrapyfi111Wrapyfi is available at https://github.com/fabawi/wrapyfi (illustrated in Figure 1) framework, a Python wrapper supporting multiple middleware bindings. Wrapyfi is a simpler alternative to GenoM3 (Mallet et al., 2010). GenoM3 adopts a model-driven approach and uses templates to define the components and data exchanges across middleware. Since it is specifically developed for Python scripting, Wrapyfi eliminates the need for having to learn another language or to define templates, unlike GenoM3. REMS (Tanaka and Mehta, 2022) is a middleware built in Python with simplistic interfaces for educational purposes. Although REMS supports a large set of robots and simulation environments, it does not address interoperability between different middleware operating on them.
@@ 3. Communication schemes
Forwarding. The forwarding scheme in Wrapyfi enables passing arguments to multiple methods, each with a different middleware setting. This forms a chain of methods, transferring arguments and return values across middleware and topics. Forwarding employs multiple scripts with unique functions, connected by register decorators, making it suitable for creating multi-step processes with several scripts having partial component support. In LABEL:fig:forwarding_example, we demonstrate data transmission between a system without ZeroMQ support and another without Yarp support, using an intermediary system that supports both. The first system dispatches the message using Yarp by invoking send_yarp. The intermediary system then forwards it using ZeroMQ to send_zmq. The final system, with Yarp disabled, receives the message via ZeroMQ by listening to send_zmq. This scheme is needed when having to follow strict specifications regarding software and middleware compatibility across systems, such as in the case of robots.

### 19 Jan 2024 v5

@@ Abstract.
http://software.knowledge-technology.info#wrapyfi.
@@ 1. Introduction
To improve interoperability between different robotic platforms and reduce reliance on a particular middleware, we have developed the open source Wrapyfi111https://github.com/fabawi/wrapyfi (illustrated in Figure 1) framework, a Python wrapper supporting multiple middleware bindings. Wrapyfi is a simpler alternative to GenoM3 (Mallet et al., 2010). GenoM3 adopts a model-driven approach and uses templates to define the components and data exchanges across middleware. Since it is specifically developed for Python scripting, Wrapyfi eliminates the need for having to learn another language or to define templates, unlike GenoM3. REMS (Tanaka and Mehta, 2022) is a middleware built in Python with simplistic interfaces for educational purposes. Although REMS supports a large set of robots and simulation environments, it does not address interoperability between different middleware operating on them.
@@ 3. Communication schemes
Forwarding. The forwarding scheme in Wrapyfi enables passing arguments to multiple methods, each with a different middleware setting. This forms a chain of methods, transferring arguments and return values across middleware and topics. Forwarding employs multiple scripts with unique functions, connected by register decorators, making it suitable for creating multi-step processes with several scripts having partial component support. In LABEL:fig:forwarding_example, we demonstrate data transmission between a system without ZeroMQ support and another without Yarp support, using an intermediary system that supports both. The first system dispatches the message using Yarp by invoking send_yarp. The intermediary system then forwards it using ZeroMQ to send_zmq. The final system, with Yarp disabled, receives the message via ZeroMQ by listening to send_zmq. This scheme is needed when strict specifications are required regarding compatibility of software and middleware between systems, as in the case of robots.